# Spatial-Spectral Morphological Mamba for Hyperspectral Image Classification

### v1

@@ Abstract
In recent years, Transformers have garnered significant attention for Hyperspectral Image Classification (HSIC) due to their self-attention mechanism, which provides strong classification performance. However, these models face major challenges in computational efficiency, as their complexity increases quadratically with the sequence length. The Mamba architecture, leveraging a State Space Model, offers a more efficient alternative to Transformers. This paper introduces the Spatial-Spectral Morphological Mamba (MorpMamba) model. In the MorpMamba model, a token generation module first converts the Hyperspectral Image (HSI) patch into spatial-spectral tokens. These tokens are then processed by a morphology block, which computes structural and shape information using depthwise separable convolutional operations. The extracted information is enhanced in a feature enhancement module that adjusts the spatial and spectral tokens based on the center region of the HSI sample, allowing for effective information fusion within each block. Subsequently, the tokens are refined in a multi-head self-attention block to further improve the feature space. Finally, the combined information is fed into the state space block for classification and the creation of the ground truth map. Experiments on widely used Hyperspectral (HS) datasets demonstrate that the MorpMamba model outperforms (parametric efficiency) both CNN and Transformer models. The source code will be made publicly available at https://github.com/MHassaanButt/MorpMamba.
@@ I Introduction
Hyperspectral Image Classification (HSIC) is crucial in diverse fields such as remote sensing [1], Earth observation [2], urban planning [3], agriculture [4], forestry [5], target/object detection [6], mineral exploration [7], environmental and climate monitoring [8, 9, 10], forensic sciences [11, 12], and even food, meat, and bakery processing [13, 14, 15, 16, 17]. The versatility and applicability of HSIC in these domains stem from its ability to capture and analyze spectral information across a wide range of wavelengths, providing detailed insights that are often unattainable with traditional imaging techniques. This spectral richness enables the identification and classification of materials, substances, and conditions with a high degree of accuracy and specificity. However, the efficacy of models designed to analyze Hyperspectral (HS) data hinges critically on their ability to manage and interpret this rich, high-dimensional information effectively [18]. The challenge lies in developing robust algorithms that can handle the complex data structures inherent in Hyperspectral Images (HSIs), ensuring accurate and reliable classifications across various applications.
@@
The recent success of Deep Learning (DL) has significantly advanced HSIC [19, 20, 21, 22, 23, 24]. Convolutional Neural Networks (CNNs), in particular, have played a pivotal role in this progress by effectively capturing both spatial and spectral features, which are crucial for accurate classification [1]. CNNs excel at learning hierarchical representations from HS data, enabling the extraction of intricate patterns that traditional methods often miss [25]. However, CNNs have inherent limitations, such as their reliance on local receptive fields, which may not fully exploit the global contextual information in HSIs [26]. This local focus can lead to suboptimal performance, especially in scenarios where understanding the broader spatial context is essential for accurate classification [18]. Moreover, CNNs typically require substantial amounts of labeled training data to achieve high performance, which can be a significant drawback given the scarcity of annotated HS datasets [20]. As a result, while CNNs have undeniably advanced the field of HSIC, their limitations highlight the need for more advanced architectures that can overcome these challenges and further enhance classification accuracy.
@@
In recent years, Transformer architecture has emerged as a leading approach for learning long-range contextual dependencies [27]. The inclusion of Transformer-based architectures in HSIC has led to significant advancements, often outperforming Traditional DL (TDL) models. With growing interest in Transformer models to further enhance HSIC, the field has witnessed notable improvements over TDL methods [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Unlike CNNs, which are limited by their local receptive fields, Transformer models excel at capturing long-range dependencies and global contextual information. This capability is particularly beneficial for HSIs, which contain rich spectral information across numerous bands. By leveraging self-attention mechanisms, Transformers can effectively model the complex relationships between different spectral bands and spatial regions, leading to more accurate and robust classification results [25]. However, despite these advancements, Transformers face substantial challenges related to computational efficiency, as their complexity scales quadratically with sequence length. This scaling issue is especially problematic in HSIC, where the high dimensionality of HS data requires significant computational resources. Consequently, the inefficiency of Transformers in managing large datasets and their reliance on extensive labeled data limit their practical application in this field.
@@
The Mamba architecture, with its State Space Model (SSM) inspired by control theory, presents a more efficient alternative to the Transformer architecture [39]. By replacing the attention mechanism with SSMs, Mamba achieves linear complexity scaling with sequence length, drastically reducing the computational overhead [39, 40]. This efficiency is particularly beneficial for HSIC, where handling long spectral sequences is crucial. Mamba’s ability to manage long sequences more efficiently, combined with its hardware-aware optimizations, makes it a superior choice for processing the vast amounts of data inherent in HSIs.
@@
Yao et al. [41] introduced SpectralMamba, leveraging the Mamba architecture for HSIC. Their approach integrates the Mamba framework with DL techniques to address inefficiencies common in sequential architectures like Recurrent Neural Networks (RNNs) and Transformers. SpectralMamba features a Gated Spatial-Spectral Merging (GSSM) module and a Piece-wise Sequential Scanning (PSS) strategy, effectively encoding spatial regularity and spectral uniqueness, which enhances discriminative representation while ensuring computational efficiency. However, the authors acknowledge limitations, including opportunities to reduce computational overhead and improve handling of spectral variability. Huang et al. [42] proposed the Spatial-Spectral Mamba (SSMamba) architecture for HSIC. This model capitalizes on the Mamba framework’s efficiency and long-range feature extraction capabilities by incorporating a spectral-spatial token generation module and multiple stacked spectral-spatial Mamba blocks. Each block consists of two basic Mamba blocks and a feature enhancement module to facilitate the fusion of spectral and spatial information. Despite its innovative design, the authors recognized the need for further optimization to robustly manage extremely high-dimensional data and explore strategies for better generalization across diverse HS datasets.
@@
Wang et al. [43] introduced the $S^{2}$amba architecture for HSIC, which integrates spatial and spectral features using SSMs. This model employs Patch Cross Scanning and Bi-directional Spectral Scanning to capture spatial and spectral features, respectively, merging them with a Spatial-Spectral Mixture Gate to enhance classification accuracy. Nonetheless, the authors noted the necessity for further optimization to effectively handle complex spatial-spectral interactions and improve computational efficiency. He et al. [44] developed the 3D-Spectral-Spatial Mamba (3DSSMamba) architecture for HSIC. This framework utilizes the strengths of the Mamba architecture by incorporating a Spectral-Spatial Token Generation module (SSTG) and a 3D Spectral-Spatial Selective Scanning (3DSS) mechanism, enabling the model to effectively capture global spectral-spatial contextual dependencies while maintaining linear computational complexity. Despite improved classification performance, the authors stressed the need for further optimization to robustly handle high-dimensional data and explore additional strategies for better generalization across diverse HS datasets.
@@
Sheng et al. [45] introduced DualMamba, a spatial-spectral Mamba architecture for HSIC. This design integrates Mamba with CNNs, effectively capturing complex spectral-spatial relationships while ensuring computational efficiency. However, the authors noted limitations, including redundancy in multi-directional scanning strategies and challenges in fully leveraging spectral information. Similarly, Zhou et al. [46] developed another spatial-spectral Mamba architecture, featuring a centralized Mamba-Cross-Scan mechanism that transforms HSI data into diverse sequences, enhancing feature extraction through a Tokenized Mamba encoder. Despite its strengths, this method is sensitive to variations in peripheral pixels and requires significant computational resources for larger patches. Yang et al. [47] proposed GraphMamba, which enhances spatial-spectral feature extraction through components like HyperMamba and SpatialGCN, addressing efficiency and contextual awareness. Nevertheless, the authors identified shortcomings, such as the need for optimization of encoding modules to accommodate diverse HS datasets and potential overfitting in high-dimensional spaces.
Therefore, this paper introduces a Morphology-based Spatial-Spectral Mamba (MorpMamba) architecture to address the limitations of previous HSIC models. Firstly: Erosion emphasizes the boundaries and edges of objects, which is crucial for distinguishing between different regions in HSIs. Dilation, on the other hand, fills gaps and connects disjointed parts of objects, enhancing the structural integrity of features. Both erosion and dilation are effective in reducing noise by smoothing out small irregularities and highlighting prominent features. By integrating these operations into the tokenization process, MorpMamba ensures that the spatial and spectral tokens capture both fine details and global structures, resulting in more robust and informative token representations. Secondly: By incorporating morphological operations in the tokenization enhancement module, our model refines the token representations, making them less sensitive to noise and artifacts. This leads to more stable and reliable features. Combined with multi-head self-attention, the model can simultaneously focus on different aspects of the spatial and spectral features, which helps capture complex relationships and dependencies within the data. Lastly: The SSM module, which models the temporal evolution of features, benefits from the enhanced and more interpretable features extracted through morphological operations. This results in a clearer understanding of how the features evolve and contribute to the final classification.
@@
Figure 1: A joint spatial-spectral feature token is first computed from the HSI using morphological operations. These tokens are then integrated into the MorpMamba model, which includes Erosion and Dilation operations, token enhancement, and a multi-head attention module. This method allows for a more selective and effective representation of information compared to standard fixed-dimension encodings. The output is then processed through a SSM, followed by normalization and a linear layer with $l_{2}^{2}$ regularization. Finally, this output is passed to the classification head for generating the ground truth.
@@
In summary, erosion and dilation help capture fine details and global structures, leading to robust and informative spatial-spectral token representations. These operations smooth out irregularities and enhance prominent features, producing stable and reliable tokens. By separately processing the spatial and spectral dimensions, the multi-head self-attention mechanism can effectively capture complex relationships and dependencies. Integrating these benefits, the proposed morphology-based tokenization process significantly improves the performance of the Mamba model for HSIC.
@@ II Proposed Methodology
where $\mathcal{X}\in\mathcal{R}^{N\times(P\times P\times C)}$, e.g., $(N\times 8\times 8\times C)$ used as input to subsequent modules, where $N$ be the total number of patches. These patches are divided into spectral and spatial patches, individually, each undergoing separate processing to generate spectral and spatial morphological profiles (Erosion and Dilation) [48]. Figure 1 provides an overview of the MorpMamba model.
@@ II-A Token Generation with Erosion and Dilation
Initially, the Erosion is performed on each spatial and spectral patch individually as Erosion: The erosion of an input tensor $\mathcal{X}$ by a kernel $\mathbf{k}$ is defined as:
@@
where the addition $\mathcal{X}(\mathbf{i})~{}\boxplus~{}\mathbf{k}(\mathbf{i}-\mathbf{j})$ is performed element-wise, where $\boxplus$ represent the dilations operation. A dilated patch is produced by combining the input patch with the Structuring Element (SE), selecting the pixel with the maximum value in the local neighborhood. As a result of the dilation procedure, the boundaries of the foreground objects in the input patch are broadened. In other words, the SE kernel’s size affects the apparent texture size for various regions within the patch and token. The Erosion (equation 2) and Dilation (equation 3) operations are performed per channel using depthwise 2D convolution layers with kernel size $(5\times 5)$ and the same pending, initialized with weights set to one (representing the SE) and the sign is inverted.
The spatial-spectral morphology combines the spatial and spectral aspects of the input tensor $\mathcal{X}$ to generate spatial-spectral tokens. The spatial morphology is obtained by applying the erosion and dilation operations to the input tensor $\mathcal{X}$ along the spatial dimensions (H and W):
@@
The resulting spatial features are then concatenated and passed through a $1\times 1$ convolution layer to obtain the combined spatial token:
$\mathbf{t}_{\text{spatial}}=\text{DepthwiseConv2D}(\operatorname*{\parallel}(\mathcal{X}_{\text{spatial}}^{\text{eroded}},\mathcal{X}_{\text{spatial}}^{\text{dilated}}))$ (6)
@@
The resulting spectral features are then transposed back to the original shape and concatenated, followed by a $1\times 1$ convolution layer to obtain the combined spectral token:
$\mathbf{t}_{\text{spectral}}=\text{DepthwiseConv2D}(\operatorname*{\parallel}(\mathcal{X}_{\text{spectral}}^{\text{eroded}},\mathcal{X}_{\text{spectral}}^{\text{dilated}}))$ (9)
The final output of the class (EDSSTokenGeneration) layer is the pair of combined spatial and spectral tokens $(\mathbf{t}_{\text{spatial}},\mathbf{t}_{\text{spectral}})$. The spatial-spectral morphology aims to capture both the spatial structure and the spectral (channel-wise) characteristics of the input tensor $\mathcal{X}$, providing a more comprehensive representation of the data.

### v2

@@ Abstract
In recent years, the emergence of Transformers with self-attention mechanism has revolutionized the hyperspectral image (HSI) classification. However, these models face major challenges in computational efficiency, as their complexity increases quadratically with the sequence length. The Mamba architecture, leveraging a state space model (SSM), offers a more efficient alternative to Transformers. This paper introduces the Spatial-Spectral Morphological Mamba (MorpMamba) model in which, a token generation module first converts the HSI patch into spatial-spectral tokens. These tokens are then processed by morphological operations, which compute structural and shape information using depthwise separable convolutional operations. The extracted information is enhanced in a feature enhancement module that adjusts the spatial and spectral tokens based on the center region of the HSI sample, allowing for effective information fusion within each block. Subsequently, the tokens are refined through a multi-head self-attention which further improves the feature space. Finally, the combined information is fed into the state space block for classification and the creation of the ground truth map. Experiments on widely used HSI datasets demonstrate that the MorpMamba model outperforms (parametric efficiency) both CNN and Transformer models. The source code will be made publicly available at https://github.com/MHassaanButt/MorpMamba.
@@ I Introduction
Hyperspectral image (HSI) classification is crucial in diverse fields such as remote sensing [1], Earth observation [2], urban planning [3], agriculture [4], forestry [5], target/object detection [6], mineral exploration [7], environmental and climate monitoring [8, 9], forensic sciences [10, 11], and even food, meat, and bakery processing [12, 13, 14, 15, 16]. The versatility and applicability of HSI classification in these domains stem from its ability to capture and analyze spectral information across a wide range of wavelengths, providing detailed insights often unattainable with traditional imaging techniques. This spectral richness enables the identification and classification of materials, substances, and conditions with high accuracy and specificity. However, the efficacy of models designed to analyze HSI data hinges critically on their ability to manage and interpret this rich, high-dimensional information effectively [17]. The challenge lies in developing robust algorithms to handle the complex data structures inherent in Hyperspectral Images (HSIs), ensuring accurate and reliable classifications across various applications.
@@
The recent success of Deep Learning (DL) has significantly advanced HSI classification [18, 19, 20, 21, 22, 23, 24]. Convolutional Neural Networks (CNNs), in particular, have played a pivotal role in this progress by effectively capturing both spatial and spectral features, which are crucial for accurate classification [25]. CNNs excel at learning hierarchical representations from HSI data, enabling the extraction of intricate patterns that traditional methods often miss [26]. However, CNNs have inherent limitations, such as their reliance on local receptive fields, which may not fully exploit the global contextual information in HSIs [27]. This local focus can lead to suboptimal performance, especially in scenarios where understanding the broader spatial context is essential for accurate classification [17]. Moreover, CNNs typically require substantial amounts of labeled training data to achieve high performance, which can be a significant drawback given the scarcity of annotated HSI datasets [19]. As a result, while CNNs have undeniably advanced the field of HSI classification, their limitations highlight the need for more advanced architectures that can overcome these challenges and further enhance classification accuracy.
@@
In recent years, Transformer architecture has emerged as a leading approach for learning long-range contextual dependencies [28]. The inclusion of Transformer-based architectures in HSI classification has led to significant advancements, often outperforming Traditional DL (TDL) models. With growing interest in Transformer models to further enhance HSI classification, the field has witnessed notable improvements over TDL methods [29, 30, 31, 32, 33, 34, 35, 36]. Unlike CNNs, which are limited by their local receptive fields, Transformer models excel at capturing long-range dependencies and global contextual information. This capability is particularly beneficial for HSIs, which contain rich spectral information across numerous bands. By leveraging self-attention mechanisms, Transformers can effectively model the complex relationships between different spectral bands and spatial regions, leading to more accurate and robust classification results [26]. However, despite these advancements, Transformers face substantial challenges related to computational efficiency, as their complexity scales quadratically with sequence length. This scaling issue is especially problematic in HSI classification, where the high dimensionality of HSI data requires significant computational resources. Consequently, the inefficiency of Transformers in managing large datasets and their reliance on extensive labeled data limit their practical application in this field.
@@
The Mamba architecture, with its State Space Model (SSM) inspired by control theory, presents a more efficient alternative to the Transformer architecture [37]. By replacing the attention mechanism with SSMs, Mamba achieves linear complexity scaling with sequence length, drastically reducing the computational overhead [37, 38]. This efficiency is particularly beneficial for HSI classification, where handling long spectral sequences is crucial. Mamba’s ability to manage long sequences more efficiently, combined with its hardware-aware optimizations, makes it a superior choice for processing the vast amounts of data inherent in HSIs.
@@
Yao et al. [39] for the first time leveraged the Mamba architecture in remote sensing for HSIC, and proposed SpectralMamba. Their approach integrates the Mamba framework with DL techniques to address inefficiencies common in sequential architectures like Recurrent Neural Networks (RNNs) and Transformers. SpectralMamba features a Gated Spatial-Spectral Merging (GSSM) module and a Piece-wise Sequential Scanning (PSS) strategy, effectively encoding spatial regularity and spectral uniqueness, which enhances discriminative representation while ensuring computational efficiency. However, the authors acknowledge limitations, including opportunities to reduce computational overhead and improve handling of spectral variability. Huang et al. [40] proposed the Spatial-Spectral Mamba (SSMamba) architecture for HSI classification. This model capitalizes on the Mamba framework’s efficiency and long-range feature extraction capabilities by incorporating a spectral-spatial token generation module and multiple stacked spectral-spatial Mamba blocks. Each block consists of two basic Mamba blocks and a feature enhancement module to facilitate the fusion of spectral and spatial information. Despite its innovative design, the authors recognized the need for further optimization to robustly manage extremely high-dimensional data and explore strategies for better generalization across diverse HSI datasets.
@@
Wang et al. [41] introduced the $S^{2}$amba architecture for HSI classification, which integrates spatial and spectral features using SSMs. This model employs Patch Cross Scanning and Bi-directional Spectral Scanning to capture spatial and spectral features, respectively, merging them with a Spatial-Spectral Mixture Gate to enhance classification accuracy. Nonetheless, the authors noted the necessity for further optimization to effectively handle complex spatial-spectral interactions and improve computational efficiency. He et al. [42] developed the 3D-Spectral-Spatial Mamba (3DSSMamba) architecture for HSI classification. This framework utilizes the strengths of the Mamba architecture by incorporating a Spectral-Spatial Token Generation module (SSTG) and a 3D Spectral-Spatial Selective Scanning (3DSS) mechanism, enabling the model to effectively capture global spectral-spatial contextual dependencies while maintaining linear computational complexity. Despite improved classification performance, the authors stressed the need for further optimization to robustly handle high-dimensional data and explore additional strategies for better generalization across diverse HSI datasets.
@@
Sheng et al. [43] introduced DualMamba, a spatial-spectral Mamba architecture for HSI classification. This design integrates Mamba with CNNs, effectively capturing complex spectral-spatial relationships while ensuring computational efficiency. However, the authors noted limitations, including redundancy in multi-directional scanning strategies and challenges in fully leveraging spectral information. Similarly, Zhou et al. [44] developed another spatial-spectral Mamba architecture, featuring a centralized Mamba-Cross-Scan mechanism that transforms HSI data into diverse sequences, enhancing feature extraction through a Tokenized Mamba encoder. Despite its strengths, this method is sensitive to variations in peripheral pixels and requires significant computational resources for larger patches. Yang et al. [45] proposed GraphMamba, which enhances spatial-spectral feature extraction through components like HyperMamba and SpatialGCN, addressing efficiency and contextual awareness. Nevertheless, the authors identified shortcomings, such as optimizing encoding modules to accommodate diverse HSI datasets and potential overfitting in high-dimensional spaces.
@@
Figure 1: A joint spatial-spectral feature token is first computed from the HSI using morphological operations. These tokens are then integrated into the MorpMamba model, which includes Erosion and Dilation operations, token enhancement, and a multi-head attention module. This method allows a more selective and effective representation of information as compared to standard fixed-dimension encodings. The output is then processed through an SSM, followed by feature normalization and a linear layer with $l_{2}^{2}$ regularization. Finally, this output is passed to the classification head for generating the ground truth.
Motivated by the ability of nonlinear operations to better identify/capture the hidden patterns and internal relationships between various shapes of objects found in complex remote sensing data [46]. This paper introduces a Morphology-based Spatial-Spectral Mamba (MorpMamba) architecture to address the limitations of previous HSI classification models. Firstly: Erosion emphasizes the boundaries and edges of objects, which is crucial for distinguishing between different regions in HSIs. Dilation, however, fills gaps and connects disjointed parts of objects, enhancing the structural integrity of features. Both erosion and dilation are effective in reducing noise by smoothing out small irregularities and highlighting prominent features. By integrating these operations into the tokenization process, MorpMamba ensures that the spatial and spectral tokens capture both fine details and global structures, resulting in more robust and informative token representations. Secondly: By incorporating morphological operations in the tokenization enhancement module, our model refines the token representations, making them less sensitive to noise and artifacts. This leads to more stable and reliable features. Combined with multi-head self-attention, the model can simultaneously focus on different aspects of the spatial and spectral features, which helps capture complex relationships and dependencies within the data. Lastly: The SSM module, which models the temporal evolution of features, benefits from the enhanced and more interpretable features extracted through morphological operations. This results in a clearer understanding of how the features evolve and contribute to the final classification.
@@
In summary, erosion and dilation help capture fine details and global structures, leading to robust and informative spatial-spectral token representations. These operations smooth out irregularities and enhance prominent features, producing stable and reliable tokens. By separately processing the spatial and spectral dimensions, the multi-head self-attention mechanism can effectively capture complex relationships and dependencies. Integrating these benefits, the proposed morphology-based tokenization process significantly improves the performance of the Mamba model for HSI classification.
@@ II Proposed Methodology
where $\mathcal{X}\in\mathcal{R}^{N\times(P\times P\times C)}$, e.g., $(N\times 8\times 8\times C)$ used as input to subsequent modules, where $N$ be the total number of 3D patches. These patches are divided into spectral and spatial sub-patches, individually, each undergoing separate processing to generate spectral and spatial morphological profiles (Erosion and Dilation) [47]. Figure 1 provides an overview of the MorpMamba architecture.
@@ II-A Token Generation with Erosion and Dilation
Initially, the erosion operation is performed on each spatial and spectral patch individually and the erosion of an input tensor $\mathcal{X}$ by a kernel $\mathbf{k}$ is defined as:
@@
where the addition $\mathcal{X}(\mathbf{i})~{}\boxplus~{}\mathbf{k}(\mathbf{i}-\mathbf{j})$ is performed element-wise, where $\boxplus$ represent the dilations operation. A dilated patch is produced by combining the input patch with the structuring element (SE), selecting the pixel with the maximum value in the local neighborhood. As a result of the dilation operation, the boundaries of the foreground objects in the input patch are broadened. In other words, the SE kernel’s size affects the apparent texture size for various regions within the patch and token. The erosion (Equation 2) and dilation (Equation 3) operations are performed per channel using depthwise 2D convolution layers with a kernel of size $(5\times 5)$ and the same pending, initialized with weights set to one (representing the SE) and the sign is inverted.
The spatial-spectral morphology combines the spatial and spectral aspects of the input tensor $\mathcal{X}$ to generate spatial-spectral tokens. The spatial morphological feature is obtained by applying the erosion and dilation operations to the input tensor $\mathcal{X}$ along the spatial dimensions ($H$ and $W$):
@@
where the resulting spatial features are then concatenated and passed through a depthwise-convolution layer to obtain the combined spatial token:
$\mathbf{t}_{\text{spatial}}=\text{Conv2D${}_{\text{Depthwise}}$}(\operatorname*{\parallel}(\mathcal{X}_{\text{spatial}}^{\text{eroded}},\mathcal{X}_{\text{spatial}}^{\text{dilated}}),\mathcal{W})$ (6)
@@
where the resulting spectral features are then concatenated and passed through a depthwise-convolution layer to obtain the combined spectral token:
$\mathbf{t}_{\text{spectral}}=\text{Conv2D${}_{\text{Depthwise}}$}(\operatorname*{\parallel}(\mathcal{X}_{\text{spectral}}^{\text{eroded}},\mathcal{X}_{\text{spectral}}^{\text{dilated}}),\mathcal{W})$ (9)
where $\mathbf{t}_{\text{spectral}}\in\mathbb{R}^{N\times(P\times P\times C)}$ as the processing is similar to spatial tokens but focuses on spectral dimensions. The final output of the class (EDSSTokenGeneration) layer is the pair of combined spatial and spectral tokens $(\mathbf{t}_{\text{spatial}},\mathbf{t}_{\text{spectral}})$. The spatial-spectral morphology aims to capture both the spatial structure and the spectral (channel-wise) characteristics of the input tensor $\mathcal{X}$, providing a more comprehensive representation of the input data.